The product of two divergent sequences is divergent. True or false? Justify.

1
Expert's answer
2022-03-15T16:09:11-0400

False


Let the two sequences be X and Y:

X = (1, 0, 2, 0, 3, 0, ............. )

Y = (0, 1, 0, 2, 0, 3, ..............)

clearly, both sequences are unbounded and diverging to "\infty"

there product X"\cdot"Y = (0, 0, 0, 0, 0, 0, .........)

which is a constant sequence converging to 0

therefore,

X"\cdot"Y is a convergent sequence and not a divergent sequence